nightshade0773:

help d-4=7/10 solve for d

FexrlessRxby:

So this question is easy. First you have to get a common denometer for the 4 and 7/10. Then you subtract whatever the common denometer between 4 and 7/10 and that will equal D.

nightshade0773:

really i did not know that tbh

FexrlessRxby:

@nightshade0773 wrote:
really i did not know that tbh
Are you pulling my leg, or are you being serious right now?

nightshade0773:

im being serious

FexrlessRxby:

@nightshade0773 wrote:
im being serious
Then glad I could help!

nightshade0773:

thx

supie:

\(d\color{Aqua}{-4}=\color{lime}{\frac{7}{10}} \) so first you're gonna wanna change the minus to a plus in `-4` then add 4 to both sides of the equation, \(\underset{\color{red}{\cancel{\Large{+4}}}{~~~~~~~}\color{lime}{+4}}{d\color{red}{\cancel{-4}}=\color{lime}{\frac{7}{10}}}\) so what's 7/10+4?

nightshade0773:

i got 4.7

supie:

yes, 4.7 or \(4\frac{7}{10}\) or \(\frac{47}{10}\) so, \(d=4.7\)

nightshade0773:

oh thx

supie:

you're welcome